OpenStudy (anonymous):

A=77 deg, B=42 deg, c=9 how do i set this up in the laws of sines or cosines to find length of sides i think i may have done this wrong on my previous problems.

OpenStudy (radar):

I believe you can use the law of sines a/sinA=c/sinC Angle C = 180-77-42=71 degrees

OpenStudy (radar):

a/sin 77 = 9/sin 71 degrees. That should do it. Solve for a, you will then have two sides all three angles, only have side b to figure out. Use law of sines again.